# Quench Spectroscopy for Dissipative and Non-Hermitian Quantum Lattice Models

### v1

@@ 3 Quench dynamics of the isolated Bose-Hubbard chain
$G_{2}(R,t)=\frac{2\bar{n}}{L}\sum_{k\neq 0}\cos(kR)\left[\operatorname{Re}(F_{k}(t))-\operatorname{Re}(F_{k}(0))+G_{k}(t)-G_{k}(0)\right].$ (28)
Note that $G_{1}(R,0)=G_{2}(R,0)=0$ as expected since both $G_{1}$ and $G_{2}$ refer to the connected version of the one-body and density-density correlation function respectively. The previous analytical expression of $G_{2}(R,t)$ is consistent with Ref. [59] where the following expression is found:
$g_{2}(R,t)=n(t)^{2}+\frac{2n(t)}{L}\sum_{k\neq 0}\cos(kR)\left[\operatorname{Re}(F_{k}(t))+G_{k}(t)\right],$ (29)
where $g_{2}(R,t)=\langle\hat{n}_{R}\hat{n}_{0}\rangle_{t}$ denotes the non-connected density-density correlation function in the SF-mean-field regime while considering loss processes via a time-dependent density $n(t)$. The latter expression also relies on the mean-field approximation. By considering particle-number conservation, i.e. $n(t)=n=\bar{n}$, for the isolated BH chain and according to the definition of $G_{2}$ and $g_{2}$, we can deduce that $G_{2}(R,t)=g_{2}(R,t)-g2(R,0)$. Hence, we find:
$G_{2}(R,t)=g_{2}(R,t)-g2(R,0)=\bar{n}^{2}+\frac{2\bar{n}}{L}\sum_{k\neq 0}\cos(kR)\left[\operatorname{Re}(F_{k}(t))+G_{k}(t)\right]-\bar{n}^{2}-\frac{2\bar{n}}{L}\sum_{k\neq 0}\cos(kR)\left[\operatorname{Re}(F_{k}(0))+G_{k}(0)\right],$ (30)
leading finally to the expected expression of $G_{2}$ at Eq. (28). Note that the connected part of the $G_{2}$ correlation function namely $-\langle\hat{n}_{R}\rangle_{t}\langle\hat{n}_{0}\rangle_{t}+\langle\hat{n}_{R}\rangle_{0}\langle\hat{n}_{0}\rangle_{0}$ vanishes for the isolated BH chain. Indeed, according to Ref. [59], in the presence of particle losses the $G_{1}$ (non-connected) one-body correlation function also called one-body density matrix is defined and takes the following form in the SF-mean-field regime:
@@
where the mean-field approximation has been considered to deduce the latter form being spatial-independent. Using the translational invariance of the model, we end up with $\langle\hat{n}_{0}\rangle_{t}=\langle\hat{n}_{R}\rangle_{t}$, $\forall R\in\mathbb{N}$. Finally, for the isolated BH chain, we have $\langle\hat{n}_{0}\rangle_{t}=\langle\hat{n}_{R}\rangle_{t}=\bar{n}=\langle\hat{n}_{0}\rangle_{0}=\langle\hat{n}_{R}\rangle_{0}$ and thus we recover that the connected part of $G_{2}$ vanishes as expected.
@@
The $G_{2}$ equal-time connected density-density correlation function at Eq. (28) is deduced theoretically by injecting the quadratic bosonic correlators $F$ and $G$ by their values obtained by solving their respective EoM. The second theoretical approach is given by the QP theory, see Ref. [68], leading to:
$G_{2}(R,t)=\frac{1}{L}\sum_{k\neq 0}\mathcal{S}^{(2)}_{k}\cos(kR)\sin^{2}(\mathcal{E}_{k,\mathrm{f}}t),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{S}^{(2)}_{k}=\frac{2\bar{n}\left(\mathcal{A}_{k,\mathrm{f}}\mathcal{B}_{k,\mathrm{i}}-\mathcal{A}_{k,\mathrm{i}}\mathcal{B}_{k,\mathrm{f}}\right)}{\left(\mathcal{A}_{k,\mathrm{f}}+\mathcal{B}_{k,\mathrm{f}}\right)\mathcal{E}_{k,\mathrm{i}}}.$ (32)
@@
On Fig. 2, both theoretical approaches are compared and we find a perfect agreement between them. This validates for another observable the EoM approach defined by the set of coupled differential equations at Eq. (16). Note that the set of EoMs at Eq. (19) is valid for the Hamiltonian $\hat{H}$ associated to the BH chain. In the SF-mean-field regime, the latter can be expressed in the generic quadratic Bose form when considering the reciprocal space:
$\hat{H}=\frac{1}{2}\sum_{k\neq 0}\mathcal{A}_{k}\left(\hat{b}^{{\dagger}}_{k}\hat{b}_{k}+\hat{b}_{-k}\hat{b}^{{\dagger}}_{-k}\right)+\mathcal{B}_{k}\left(\hat{b}^{{\dagger}}_{k}\hat{b}^{{\dagger}}_{-k}+\hat{b}_{k}\hat{b}_{-k}\right)=\frac{1}{2}\sum_{k\neq 0}\begin{pmatrix}\hat{b}^{{\dagger}}_{k}&\hat{b}_{-k}\end{pmatrix}\hat{H}_{\mathrm{BdG}}(k)\begin{pmatrix}\hat{b}_{k}\\ \hat{b}^{{\dagger}}_{-k}\end{pmatrix},\leavevmode\nobreak\ \hat{H}_{\mathrm{BdG}}(k)=\begin{pmatrix}\mathcal{A}_{k}&\mathcal{B}_{k}\\ \mathcal{B}_{k}&\mathcal{A}_{k}\end{pmatrix},$ (33)
with $\mathcal{A}_{k}=4J\sin^{2}(k/2)+U\bar{n}$ and $\mathcal{B}_{k}=U\bar{n}$ [69, 58] and $\hat{H}_{\mathrm{BdG}}$ corresponding to the bosonic Bogolyubov-de-Gennes Hamiltonian. Similar statements are expected to apply for Hamiltonians which can be cast into the previous form and involving a fermionic Bogolyubov-de-Gennes Hamiltonian.
@@
By considering the thermodynamic limit, i.e. $L\rightarrow+\infty$, and by performing some trivial algebra, for both connected equal-time correlation functions can be expressed in the following generic form  [57]:
$G_{\mathrm{i}}(R,t)\sim\int_{\mathcal{B}}\mathrm{d}k\leavevmode\nobreak\ \mathcal{S}_{k}^{(\mathrm{i})}\left\{e^{i(kR+2\mathcal{E}_{k,\mathrm{f}}t)}+e^{i(kR-2\mathcal{E}_{k,\mathrm{f}}t)}\right\},$ (34)
@@
where $\mathcal{B}=[-\pi,\pi]$ denotes the first Brillouin zone. Note that the latter form unveiled at Ref. [57] will be essential for the discussion about the quench spectroscopy method (QS) applied to the dissipative Bose-Hubbard chain. More precisely, Eq. (37) consists in the generic form of equal-time connected correlation functions (ETCCF) for an isolated quantum lattice model driven out of equilibrium via a weak sudden global quench. In Ref. [57], the large distance and time behavior of the latter expression is characterized using the stationary phase approximation to unravel general statements regarding the correlation spreading in closed quantum systems. For short-range interacting quantum models, the space-time pattern of ETCCFs is expected to display a twofold linear structure. This double structure is characterised by a ballistic, i.e. linear, correlation edge (CE) separating the causal region of the correlations from the non-causal one. The causal region refers to the region where the correlations are different from zero. In other words, beyond the CE, the space-time correlations are exponentially suppressed. The second structure consists in the ballistic spreading of a series of local maxima and minima in the vicinity of the CE. The CE velocity is given by twice the maximal group velocity corresponding to the quasiparticle pair with the highest velocity, i.e. $V_{\mathrm{CE}}=2V_{\mathrm{g}}(k^{*})=2\mathrm{max}(\mathrm{d}\mathcal{E}_{k,\mathrm{f}}/\mathrm{d}k)$ with $k^{*}=\mathrm{argmax}(\mathrm{d}\mathcal{E}_{k,\mathrm{f}}/\mathrm{d}k)$. The velocity associated to the linear spreading of the local extremum is given by twice the phase velocity evaluated at the quasimomentum for which the group velocity is maximal, i.e. $V_{\mathrm{m}}=2V_{\varphi}(k^{*})=2\mathcal{E}_{k^{*}}/k^{*}$. The validity of the latter behaviour has been certified both analytically [68, 57] and numerically using whether time-dependent matrix product state calculations [58] or time-dependent variational Monte-Carlo calculations [55].
4 Double quench dynamics of the Bose-Hubbard chain
@@
In what follows, we investigate the behaviour of $G_{2}(R,t)$ for a sudden global quench not only on the dissipation strength from $\gamma=0$ to $\gamma>0$ but also on the interaction strength $U$ from $U=U_{\mathrm{i}}$ to $U=U_{\mathrm{f}}$ such that the BH chain is initially confined in the SF phase. Experimentally, the quench on $U$ can be realised in ultracold-atom experiments by suddenly lowering or raising the depth of the optical lattice in order to decrease or to increase the dimensionless interaction parameter $U/J$ respectively [3]. In this context, the EoMs at Eq. (16) as well as the corresponding initial conditions at Eq. (17d) have to be slightly modified. In the EoMs, $\mathcal{A}_{k}$ and $\mathcal{B}_{k}$ are replaced by $\mathcal{A}_{k,\mathrm{f}}$ and $\mathcal{B}_{k,\mathrm{f}}$ depending now on the post-quench interaction strength $U_{\mathrm{f}}$. For the initial conditions, $\mathcal{A}_{k}$ and $\mathcal{B}_{k}$ are replaced by $\mathcal{A}_{k,\mathrm{i}}$ and $\mathcal{B}_{k,\mathrm{i}}$ depending now on the pre-quench interaction strength $U_{\mathrm{i}}$. $G_{2}$ the ETCCF to investigate density fluctuations is computed by solving the previous set of EoMs whose values of the relevant correlators are injected in the theoretical expression of $g_{2}$ given by:
$g_{2}(R,t)=n(t)^{2}+\frac{2n(t)}{L}\sum_{k\neq 0}\cos(kR)\left[G_{k}(t)+\operatorname{Re}(F_{k}(t))\right].$ (35)
@@
Figure 3: Space-time pattern of (left) $G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{\mathrm{c}}$ (right) $G_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}-\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{0}$ for a sudden global quench on the dissipation strength from $\gamma=0$ to $\gamma>0$ and on the two-body repulsive interaction strength $U$ from $U=U_{\mathrm{i}}$ to $U=U_{\mathrm{f}}$ such that the quantum system is initially confined in the SF phase. The solid green line represents the correlation edge (CE) propagating linearly at the velocity $V_{\mathrm{CE}}=2V_{\mathrm{g}}(k^{*})=2\mathrm{max}(\mathrm{d}\mathcal{E}_{k,\mathrm{f}}/\mathrm{d}k)$. The solid blue line represents the ballistic spreading of the series of local maxima and minima in the vicinity of the CE at the velocity $V_{\mathrm{m}}=2V_{\varphi}(k^{*})=2\mathcal{E}_{k^{*},\mathrm{f}}/k^{*}$. The parameters are: $N(0)=L=400$ to satisfy the thermodynamic limit, $J=1$, $U_{\mathrm{f}}=0.5$, $U_{\mathrm{i}}=2$, $\gamma=0.1$. For the twofold linear structure induced by the quench on the interactions to be visible, the amplitudes of both patterns have been saturated.
@@
The latter is plotted on Fig. 3 and we clearly see the effect of each quench individually. The quench on $U$ gives rise to a twofold linear structure characterised by a correlation edge (CE) propagating ballistically at twice the maximal group velocity. The second structure consists in a series of local maxima and minima spreading also ballistically in the vicinity of the CE at twice the phase velocity evaluated at the quasimomentum for which the group velocity is maximal. This is reminiscent of the behaviour of $G_{2}$ for a sudden global quench on the interaction strength $U$ for the isolated BH chain confined in the SF phase. We can thus deduced from the previous findings that the algebraic decay in time of $G_{2}$, thus independent on the lattice site index $R$, corresponds to the signature of the quench on the dissipation strength $\gamma$. The latter effect is clearly visible in the background of Fig. 3. Similar statements are also valid for the $G_{1}$ ETCCF on Fig. 3 where $G_{1}$ has the following expression since $\langle\hat{n}_{0}\rangle_{t}$ is not a conserved quantity anymore (see Eq. (23)):
$G_{1}(R,t)=\frac{1}{L}\LARGE[\langle\hat{n}_{0}\rangle_{t}-\langle\hat{n}_{0}\rangle_{0}+\sum_{k\neq 0}\cos(kR)\huge(\langle\hat{n}_{k}\rangle_{t}-\langle\hat{n}_{k}\rangle_{0}\huge)\LARGE]=\frac{1}{L}\LARGE\{G_{0}(t)-G_{0}(0)+\sum_{k\neq 0}\cos(kR)\huge[G_{k}(t)-G_{k}(0)\huge]\LARGE\}.$ (36)
5 Quench spectroscopy of the double-quenched Bose-Hubbard chain

### v2

@@ II.2 Quench dynamics of the isolated Bose-Hubbard chain
$G_{2}(R,t)=\frac{2\bar{n}}{L}\sum_{k\neq 0}\cos(kR)\{\operatorname{Re}[F_{k}(t)]+G_{k}(t)$
$-\operatorname{Re}[F_{k}(0)]-G_{k}(0)\},$ (20)
Note that $G_{1}(R,0)=G_{2}(R,0)=0$ as expected since both $G_{1}$ and $G_{2}$ refer to the connected version of the one-body and density-density correlation function respectively. The previous analytical form of $G_{2}(R,t)$ is consistent with the following expression found using the set of EoMs defined at Eq. (8):
$g_{2}(R,t)=n(t)^{2}+\frac{2n(t)}{L}\sum_{k\neq 0}\cos(kR)\left\{\operatorname{Re}[F_{k}(t)]+G_{k}(t)\right\},$ (21)
where $g_{2}(R,t)=\langle\hat{n}_{R}\hat{n}_{0}\rangle_{t}$ denotes the non-connected density-density correlation function in the SF-mean-field regime while considering loss processes via a time-dependent density $n(t)$. The latter expression is discussed in Appendix A together with the theoretical expression of the condensate density $n(t)$ in the non-interacting case $U=0$. By considering particle-number conservation, i.e. $n(t)=n=\bar{n}$, for the isolated BH chain and according to the definition of $G_{2}$ and $g_{2}$, we can deduce that $G_{2}(R,t)=g_{2}(R,t)-g2(R,0)$. Hence, we find:
$G_{2}(R,t)=\bar{n}^{2}+\frac{2\bar{n}}{L}\sum_{k\neq 0}\cos(kR)\left[\operatorname{Re}(F_{k}(t))+G_{k}(t)\right]$
$-\bar{n}^{2}-\frac{2\bar{n}}{L}\sum_{k\neq 0}\cos(kR)\left[\operatorname{Re}(F_{k}(0))+G_{k}(0)\right],$ (22)
leading finally to the expected expression of $G_{2}$ at Eq. (20). Note that the connected part of the $G_{2}$ correlation function namely $-\langle\hat{n}_{R}\rangle_{t}\langle\hat{n}_{0}\rangle_{t}+\langle\hat{n}_{R}\rangle_{0}\langle\hat{n}_{0}\rangle_{0}$ vanishes for the isolated BH chain. Indeed, according to Ref. despresnew , in the presence of particle losses the $G_{1}$ (non-connected) one-body correlation function also called one-body density matrix is defined and takes the following form in the SF-mean-field regime:
@@
where the mean-field approximation has been considered to deduce the latter form being spatially independent. Using the translational invariance of the model, we end up with $\langle\hat{n}_{0}\rangle_{t}=\langle\hat{n}_{R}\rangle_{t}$, $\forall R\in\mathbb{N}$. Finally, for the isolated BH chain, we have $\langle\hat{n}_{0}\rangle_{t}=\langle\hat{n}_{R}\rangle_{t}=\bar{n}=\langle\hat{n}_{0}\rangle_{0}=\langle\hat{n}_{R}\rangle_{0}$ and thus we recover that the connected part of $G_{2}$ vanishes as expected.
@@
$G_{2}$ the equal-time connected density-density correlation function at Eq. (20) is deduced theoretically by injecting the quadratic bosonic correlators $F$ and $G$ by their values obtained by solving their respective EoM. The second theoretical approach is given by the QP theory, see Ref. despres_these , leading to:
$G_{2}(R,t)=$ $\leavevmode\nobreak\ \frac{1}{L}\sum_{k\neq 0}\mathcal{S}^{(2)}_{k}\cos(kR)\sin^{2}(\mathcal{E}_{k,\mathrm{f}}t),$ (24a)
$\mathcal{S}^{(2)}_{k}=$ $\leavevmode\nobreak\ \frac{2\bar{n}\left(\mathcal{A}_{k,\mathrm{f}}\mathcal{B}_{k,\mathrm{i}}-\mathcal{A}_{k,\mathrm{i}}\mathcal{B}_{k,\mathrm{f}}\right)}{\left(\mathcal{A}_{k,\mathrm{f}}+\mathcal{B}_{k,\mathrm{f}}\right)\mathcal{E}_{k,\mathrm{i}}}.$ (24b)
@@
On Fig. 2, both theoretical approaches are compared and we find a perfect agreement between them. This validates for another observable the EoM approach defined by the set of coupled differential equations at Eq. (8). Note that the set of EoMs at Eq. (11) is valid when considering the Hamiltonian $\hat{H}$ of the BH chain confined in the SF-mean-field regime which can be expressed in a generic quadratic Bose form in reciprocal space:
$\hat{H}=\frac{1}{2}\sum_{k\neq 0}\mathcal{A}_{k}\left(\hat{b}^{{\dagger}}_{k}\hat{b}_{k}+\hat{b}_{-k}\hat{b}^{{\dagger}}_{-k}\right)+\mathcal{B}_{k}\left(\hat{b}^{{\dagger}}_{k}\hat{b}^{{\dagger}}_{-k}+\hat{b}_{k}\hat{b}_{-k}\right),$ (25)
with $\mathcal{A}_{k}=4J\sin^{2}(k/2)+U\bar{n}$ and $\mathcal{B}_{k}=U\bar{n}$, see Eq. (3c) and Refs. roux2013 ; despres2019 . The latter can be reformulated in a Bogolyubov-de-Gennes form and reads as:
$\hat{H}=\frac{1}{2}\sum_{k\neq 0}\begin{pmatrix}\hat{b}^{{\dagger}}_{k}&\hat{b}_{-k}\end{pmatrix}\hat{H}_{\mathrm{BdG}}(k)\begin{pmatrix}\hat{b}_{k}\\ \hat{b}^{{\dagger}}_{-k}\end{pmatrix};$ (26)
$\hat{H}_{\mathrm{BdG}}(k)=\begin{pmatrix}\mathcal{A}_{k}&\mathcal{B}_{k}\\ \mathcal{B}_{k}&\mathcal{A}_{k}\end{pmatrix},$ (27)
where $\hat{H}_{\mathrm{BdG}}$ refers to the bosonic Bogolyubov-de-Gennes Hamiltonian. Similar statements will apply for any quantum system whose Hamiltonian can be reformulated in a quadratic Bose form, i.e. can be cast into the previous form. The latter statements are also valid for fermionic quantum systems provided that their Hamiltonian displays a quadratic Fermi form, i.e. can be expressed via a fermionic Bogolyubov-de-Gennes Hamiltonian. In Section 8, the latter expectations are discussed by considering the $s=1/2$ short-range interacting transverse-field Ising chain in the paramagnetic phase.
@@
By considering the thermodynamic limit, i.e. $L\rightarrow+\infty$, and by performing some trivial algebra, for both connected equal-time correlation functions can be expressed in the following generic form  despres2018 :
$G_{\mathrm{i}}(R,t)\sim\int_{\mathcal{B}}\mathrm{d}k\mathcal{S}_{k}^{(\mathrm{i})}\left[e^{i(kR+2\mathcal{E}_{k,\mathrm{f}}t)}+e^{i(kR-2\mathcal{E}_{k,\mathrm{f}}t)}\right],$ (28)
@@
where $\mathcal{B}=[-\pi,\pi]$ denotes the first Brillouin zone. Note that the latter form unveiled at Ref. despres2018 will be essential for the discussion about the quench spectroscopy method (QS) applied to the dissipative Bose-Hubbard chain. More precisely, Eq. (31) consists in the generic form of equal-time connected correlation functions (ETCCF) for an isolated quantum lattice model driven out of equilibrium via a weak sudden global quench. In Ref. despres2018 , the large distance and time behavior of the latter expression is characterized using the stationary phase approximation to unravel general statements regarding the correlation spreading in closed quantum systems. For short-range interacting quantum models, the space-time pattern of ETCCFs is expected to display a twofold linear structure. This double structure is characterized by a ballistic, i.e. linear, correlation edge (CE) separating the causal region of the correlations from the non-causal one. The causal region refers to the region where the correlations are different from zero. In other words, beyond the CE, the space-time correlations are exponentially suppressed. The second structure consists in the ballistic spreading of a series of local maxima and minima in the vicinity of the CE. The CE velocity is given by twice the maximal group velocity corresponding to the quasiparticle pair with the highest velocity, i.e. $V_{\mathrm{CE}}=2V_{\mathrm{g}}(k^{*})=2\mathrm{max}(\mathrm{d}\mathcal{E}_{k,\mathrm{f}}/\mathrm{d}k)$ with $k^{*}=\mathrm{argmax}(\mathrm{d}\mathcal{E}_{k,\mathrm{f}}/\mathrm{d}k)$. The velocity associated to the linear spreading of the local extremum is given by twice the phase velocity evaluated at the quasi-momentum for which the group velocity is maximal, i.e. $V_{\mathrm{m}}=2V_{\varphi}(k^{*})=2\mathcal{E}_{k^{*}}/k^{*}$. The validity of the latter behavior has been certified both analytically despres_these ; despres2018 and numerically using whether time-dependent matrix product state calculations despres2019 or time-dependent variational Monte-Carlo calculations carleo2014 .
II.3 Double quench dynamics of the Bose-Hubbard chain
@@
In what follows, we investigate the behavior of $G_{2}(R,t)$ for a sudden global quench not only on the dissipation strength from $\gamma=0$ to $\gamma>0$ but also on the interaction strength $U$ from $U=U_{\mathrm{i}}$ to $U=U_{\mathrm{f}}$ such that the BH chain is initially confined in the SF phase. Experimentally, the quench on $U$ can be realized in ultra-cold-atom experiments by suddenly lowering or raising the depth of the optical lattice in order to decrease or to increase the dimensionless interaction parameter $U/J$ respectively cheneau2012 . In this context, the EoMs at Eq. (8) as well as the corresponding initial conditions at Eq. (9) have to be slightly modified. In the EoMs, $\mathcal{A}_{k}$ and $\mathcal{B}_{k}$ are replaced by $\mathcal{A}_{k,\mathrm{f}}$ and $\mathcal{B}_{k,\mathrm{f}}$ depending now on the post-quench interaction strength $U_{\mathrm{f}}$. For the initial conditions, $\mathcal{A}_{k}$ and $\mathcal{B}_{k}$ are replaced by $\mathcal{A}_{k,\mathrm{i}}$ and $\mathcal{B}_{k,\mathrm{i}}$ depending now on the pre-quench interaction strength $U_{\mathrm{i}}$. $G_{2}$ the ETCCF to investigate density fluctuations is computed by solving the previous set of EoMs whose values of the relevant correlators are injected in the theoretical expression of $g_{2}$ given by:
$g_{2}(R,t)=n(t)^{2}+\frac{2n(t)}{L}\sum_{k\neq 0}\cos(kR)\left\{G_{k}(t)+\operatorname{Re}[F_{k}(t)]\right\}.$ (29)
@@
Figure 3: Space-time pattern of (top panel) $G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{\mathrm{c}}$ (bottom panel) $G_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}-\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{0}$ for a sudden global quench on the dissipation strength from $\gamma=0$ to $\gamma>0$ and on the two-body repulsive interaction strength $U$ from $U=U_{\mathrm{i}}$ to $U=U_{\mathrm{f}}$ such that the quantum system is initially confined in the SF-mean-field regime. The solid green line represents the correlation edge (CE) propagating linearly at the velocity $V_{\mathrm{CE}}=2V_{\mathrm{g}}(k^{*})=2\mathrm{max}(\mathrm{d}\mathcal{E}_{k,\mathrm{f}}/\mathrm{d}k)$. The solid blue line represents the ballistic spreading of the series of local maxima and minima in the vicinity of the CE at the velocity $V_{\mathrm{m}}=2V_{\varphi}(k^{*})=2\mathcal{E}_{k^{*},\mathrm{f}}/k^{*}$. The considered post-quench and pre-quench dimensionless interaction parameters are $U_{\mathrm{f}}n(0)/J=0.5$ and $U_{\mathrm{i}}n(0)/J=2$ respectively. The dimensionless dissipation strength is given by $\gamma/J=0.1$.
@@
The latter is plotted on Fig. 3 and we clearly see the effect of each quench individually. The quench on $U$ gives rise to a twofold linear structure characterized by a correlation edge (CE) propagating ballistically at twice the maximal group velocity. The second structure consists in a series of local maxima and minima spreading also ballistically in the vicinity of the CE at twice the phase velocity evaluated at the quasi-momentum for which the group velocity is maximal. This is reminiscent of the behavior of $G_{2}$ for a sudden global quench on the interaction strength $U$ for the isolated BH chain confined in the SF-mean-field regime. We can thus deduced from the previous findings that the algebraic decay in time of $G_{2}$, thus independent on the lattice site index $R$, corresponds to the signature of the quench on the dissipation strength $\gamma$. The latter effect is clearly visible in the background of Fig. 3. Similar statements are also valid for the $G_{1}$ ETCCF as depicted on Fig. 3 and having the following expression, see also Eq. (14), where $G_{k}(0)=\langle\hat{n}_{0}\rangle_{t}$ is no longer a conserved quantity:
$G_{1}(R,t)=$ $\leavevmode\nobreak\ \frac{1}{L}\sum_{k}\cos(kR)\left[G_{k}(t)-G_{k}(0)\right],$ (30)
while relying on the notations used to define the set of EoMs at Eq. (8).
II.4 Quench spectroscopy of the double-quenched Bose-Hubbard chain